OpenStudy (anonymous):

Janet can shovel snow from her driveway in 60 minutes. Bill can do the same job in 65 minutes. How long would it take Janet and Bill to shovel the driveway if they work together?

OpenStudy (anonymous):

Not sure, but it would take less than 60 minutes right?

OpenStudy (saifoo.khan):

156/5

OpenStudy (anonymous):

Can you show the working?

OpenStudy (saifoo.khan):

No. =D

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

the work = 1 rate of janet is 1/60 rate of bill is 1/65 rate when they are working together 1/60 + 1/65 = 5/156 1/(5/156) = 156/5

OpenStudy (anonymous):

so its 156/5 minutes?

OpenStudy (anonymous):

Yes, pretty sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that simplifyed?

OpenStudy (saifoo.khan):

totally.

OpenStudy (anonymous):

kk thx